# Demystifying Spectral Feature Learning for Instrumental Variable Regression

### v1

@@ 5 Experiments
Let $d=11$, $\mathcal{X}=[0,2\pi]$, $\mathcal{Z}=[0,2\pi]$, $\sigma^{2}=0.1$, $f(x)=(\sin(x),\sin(2x),\ldots,\sin((d-1)x))$, $g(z)=(\sin(z),\sin(2z),\ldots,\sin((d-1)z))$, and let $P,Q$ be two orthogonal $(d-1)\times(d-1)$ matrices. Define $v:\mathcal{X}\to\mathbb{R}^{d-1}$ and $u:\mathcal{Z}\to\mathbb{R}^{d-1}$ as $v(x)=Pf(x)$ and $u(z)=Qg(z)$. We set $\mathcal{T}=\mathbbm{1}_{\mathcal{Z}}\otimes\mathbbm{1}_{\mathcal{X}}+\sum_{i=1}^{d-1}\sigma_{i}u_{i}\otimes v_{i}$. $(X,Z)$ is jointly sampled with rejection sampling. $Z,X$ are uniform on $[0,2\pi]$ and $X$ given $Z$ admits $\mathcal{T}$ as conditional expectation operator. We then sample $Y$ following 3. The multiplication of the trigonometric functions by the orthogonal matrices $P,Q$ ensures that the consecutive singular functions of $\mathcal{T}$ are comparably complex. By keeping the difficulty constant per singular function we can isolate the influence of the singular value decay on the performance of the final estimator. We set the singular values $\{\sigma_{i}\}_{i=1}^{d-1}$ to decay linearly, from some initial value $\sigma_{1}$ to a $\sigma_{d-1}=c_{\sigma}\sigma_{1}$, $c_{\sigma}\in[0,1]$. To control spectral alignment, we set $h_{0}=\sum_{i=1}^{d-1}\alpha_{i}v_{i}$ subject to $\|\alpha\|_{\ell_{2}}=1$, and allow $\{\alpha_{i}\}_{i=1}^{d-1}$ to decay linearly from $\alpha_{1}$ to $\alpha_{d-1}=c_{\alpha}\alpha_{1}$, $c_{\alpha}\in[0,1]$. Each value of $c_{\alpha}\in[0,1]$ controls the alignment of $h_{0}$ with the top singular functions of $\mathcal{T}$: larger values correspond to stronger alignment. Similarly, $c_{\sigma}\in[0,1]$ controls the ill-posedness of the problem: smaller values indicate faster decay of singular values and hence more severe ill-posedness.
@@
In Figure 1, we report the mean squared error (MSE) of two 2SLS estimators across a range of spectral decay rates ($c_{\sigma}$) and structural function decay rates ($c_{\alpha}$). Left: Oracle spectral features: The left panel uses 2SLS with fixed features $f$ and $g$ that form spectral features. This represents an oracle setting where feature learning is not needed. As expected, the MSE decreases from top to bottom as the problem becomes better conditioned (i.e., as $c_{\sigma}$ increases). Additionally, the MSE decreases from left to right, as smaller values of $c_{\alpha}$ correspond to stronger spectral alignment: more of $h_{0}$ is concentrated on the top singular functions of $\mathcal{T}$. These trends are fully consistent with the bound established in Corollary 1. Right: Learned features. In the right panel, we learn 50 features from data by minimizing the empirical contrastive loss (8) using a neural network. We then build a 2SLS estimator on top of the learned features. The same qualitative trends are observed: MSE decreases with slower singular value decay and better alignment. Moreover, when $c_{\sigma}$ is small (top rows), learning the entire eigenspaces becomes difficult. In this regime, performance is most sensitive to $c_{\alpha}$. In particular, if $h_{0}$ is concentrated on the top singular functions (small $c_{\alpha}$), good recovery is still possible. In contrast, when $c_{\sigma}$ is large (bottom rows), the singular functions are equally important and more easily learned, rendering the alignment of $h_{0}$ less critical. The mismatch between the oracle and learned settings is because we do not exactly recover the singular feature spaces; as predicted by Theorem 3. It is also worth noting that when $h_{0}$ is concentrated on high singular values, it may sometimes be beneficial for the singular values to decay quickly, by making it easier for the feature-learning model to learn the top singular functions well. This may explain the slight top-bottom increase in losses on the right end of the right panel.
The previous experiment assumes $h_{0}$ lies in the span of the singular functions of $\mathcal{T}$. Figure 2 illustrates three regimes. The left panel shows a well-aligned case (“good/bad”), where $h_{0}$ is entirely supported on the singular functions of $\mathcal{T}$, allowing for exact recovery as long as enough data is available. The middle panel represents a “partial alignment” regime, where $h_{0}$ has some overlap with the singular functions; in this case, recovery is imperfect, but the estimator can still capture some signal. Finally, the right panel illustrates the “misaligned” or “very ugly” regime, where $h_{0}$ is orthogonal to the singular functions. The instrument provides no information about the signal, and all spectral methods fail. We refer to Appendix C for further discussion of these scenarios and the experimental setup.
@@ A.1 Proofs of Section 2
Under Assumption 2, the joint distribution $\pi_{X,Z}$ is absolutely continuous with respect to the product measure $\pi_{X}\otimes\pi_{Z}$, and thus admits a density
@@
In order to prove 1, we use the following lemma.
@@
Under Assumption 2, for $\pi_{X}\otimes\pi_{Z}$-almost every $(x,z)$,
@@
[…]
We recall and prove 1. This is a classical result, see, e.g., [25].
@@
[…]
@@
[…]
<<$\infty$< + ∞ (by Assumption 2, since $p\in L_{2}(\pi_{X}\otimes\pi_{Z})$).
@@ A.2 Proofs of Section 3
4 is due to Lemma 1 of [3] and 2 corresponds to part (iii). We provide a full proof for completeness.
@@
We recall and prove Corollary 1:
@@
Let us verify that Assumption 3 holds under Definition 2, i.e., when $\mathcal{H}_{\varphi,d}=\mathcal{V}_{d}$ and $\mathcal{H}_{\psi,d}=\mathcal{U}_{d}$.
@@
(ii) Let $h_{0}\in L_{2}(X)$, and consider its projection onto the orthogonal complement of $\mathcal{H}_{\varphi,d}$. Since $\mathcal{T}$ is compact with singular value decomposition $\mathcal{T}=\sum_{i=1}^{\infty}\sigma_{i}u_{i}\otimes v_{i}$ under Assumption 2, we have
@@
This implies, using that $\tau_{\varphi,d}=\sigma_{d}^{-1}$ by 2,
@@
Since both parts of Assumption 3 hold, and $\tau_{\varphi,d}=\sigma_{d}^{-1}$ by 2, we may apply Theorem 1, which yields the desired result. ∎
A.3 Proofs of Section 4
The following result shows that Eq. 9 holds if and only if $\operatorname{rank}(\mathcal{T})\leq d$.
@@ A.3 Proofs of Section 4
$\Rightarrow$ Assume $p(x,z)=\varphi(x)^{\mathsf{\scriptscriptstyle T}}\psi(z)$. Then, by Eq. 10, for all $h\in L_{2}(X)$,
[…]
@@
[…]
@@
In order to prove Theorem 3 we prove intermediate results.
@@
Assumption 5 holds if and only if $\operatorname{rank}(\hat{\mathcal{T}}_{d})=d$.
@@
Assume that Assumption 5 holds, i.e., $\{\hat{\varphi}_{1},\ldots,\hat{\varphi}_{d}\}$ and $\{\hat{\psi}_{1},\ldots,\hat{\psi}_{d}\}$ are linearly independent. We show that $\operatorname{rank}(\hat{\mathcal{T}}_{d})=d$.
@@
Under Assumption 5, the family $\{\hat{\varphi}_{i}\}$ is linearly independent, so $\Phi$ is surjective. Similarly, the family $\{\hat{\psi}_{i}\}$ is linearly independent, so $A$ is injective. Therefore, the composition $A\circ\Phi$ has image of dimension
@@
Under Assumption 5, we can therefore write $\hat{\mathcal{T}}_{d}$ in the following SVD form:
@@
Under Assumption 5, $\mathcal{H}_{\hat{\varphi},d}=\operatorname{span}\left\{\hat{v}_{1},\ldots,\hat{v}_{d}\right\}$ and $\mathcal{H}_{\hat{\psi},d}=\operatorname{span}\left\{\hat{u}_{1},\ldots,\hat{u}_{d}\right\}$.
@@
We now recall and prove Theorem 3: See 3
@@
We first show that $\sigma_{d}^{-1}\leq\tau_{\hat{\varphi},d}\leq(\sigma_{d}-2\hat{\varepsilon}_{d})^{-1}$. First note that if $h\in\mathcal{H}_{\hat{\varphi},d}$, then by Proposition 7, $h\in\mathcal{V}_{d}$. Using the SVD of $\hat{\mathcal{T}}_{d}$ from Eq. 14, and writing $h=\sum_{i=1}^{d}\langle h,\hat{v}_{i}\rangle\hat{v}_{i}$, we obtain:
@@
By Wedin’s sin–$\Theta$ theorem (Theorem 2.9 and Eq. (2.26a) in [8]), and under the assumption $\hat{\varepsilon}_{d}<(1-1/\sqrt{2})\sigma_{d}$, we have:
@@
To conclude we now check Assumption 3. Let us introduce
@@
The unique use of Assumption 3-i) in the proof of Theorem B.1 [6] is to prove that $s_{\hat{\psi},\hat{\varphi},d}^{-1}=O(\tau_{\hat{\varphi},d})$ as $d\to+\infty$, which we now directly prove. First, note that for any $h\in L_{2}(X),$
@@
where we used $\left\|\mathcal{T}\right\|_{\mathrm{op}}\leq 1$, the definition of $\mathcal{T}_{d}$ and Wedin’s sin–$\Theta$ theorem. On the other hand,
@@
which is satisfied if $\hat{\varepsilon}_{d}=o(\sigma_{d}^{2})$. We therefore conclude that if $\hat{\varepsilon}_{d}=o(\sigma_{d}^{2})$ then $s_{\hat{\psi},\hat{\varphi},d}^{-1}=O(\tau_{\hat{\varphi},d})$. Finally, we check Assumption 3-ii). First note that by definition of $\Pi_{\hat{\varphi},d}$ and $\hat{\mathcal{T}}_{d}$, we have $\hat{\mathcal{T}}_{d}(\Pi_{\hat{\varphi},d})_{\perp}=0$, therefore,
@@
We now prove the equivalence between Eq. 6 and Eq. 7. While the proof strategy follows that of Kostic et al. [24, Theorem 1], we include our own version for completeness, as the parametrization of the learned operator differs. Specifically, [24] directly approximate the truncated operator $\mathcal{T}_{d}$ using a singular value decomposition of the form $\mathbbm{1}_{\mathcal{Z}}\otimes\mathbbm{1}_{\mathcal{X}}+\sum_{i=2}^{d}\hat{\sigma}_{i}\hat{\psi}_{i}\otimes\hat{\varphi}_{i}$, where the singular values $\hat{\sigma}_{i}$ are learned explicitly via a separate network. In contrast, our approach learns only the feature maps $(\varphi_{i},\psi_{i})$ and defines the approximation $\mathcal{T}_{d}(\hat{\varphi},\hat{\psi})=\sum_{i=1}^{d}\hat{\psi}_{i}\otimes\hat{\varphi}_{i}$. As such, we provide a short self-contained derivation adapted to our setting for clarity. Recall that the Eckart–Young–Mirsky formulation of the objective is defined as

### v2

@@ 5 Experiments
Let $d=11$, $\mathcal{X}=[0,2\pi]$, $\mathcal{Z}=[0,2\pi]$, $\sigma^{2}=0.1$, $f(x)=(\sin(x),\sin(2x),\ldots,\sin((d-1)x))$, $g(z)=(\sin(z),\sin(2z),\ldots,\sin((d-1)z))$, and let $P,Q$ be two orthogonal $(d-1)\times(d-1)$ matrices. Define $v:\mathcal{X}\to\mathbb{R}^{d-1}$ and $u:\mathcal{Z}\to\mathbb{R}^{d-1}$ as $v(x)=Pf(x)$ and $u(z)=Qg(z)$. We set $\mathcal{T}=\mathbbm{1}_{\mathcal{Z}}\otimes\mathbbm{1}_{\mathcal{X}}+\sum_{i=1}^{d-1}\sigma_{i}u_{i}\otimes v_{i}$. $(X,Z)$ is jointly sampled with rejection sampling. $Z,X$ are uniform on $0,2\pi]$ and $X$ given $Z$ admits $\mathcal{T}$ as conditional expectation operator. We then sample $Y$ following ˜3. The multiplication of the trigonometric functions by the orthogonal matrices $P,Q$ ensures that the consecutive singular functions of $\mathcal{T}$ are comparably complex. By keeping the difficulty constant per singular function we can isolate the influence of the singular value decay on the performance of the final estimator. We set the singular values $\{\sigma_{i}\}_{i=1}^{d-1}$ to decay linearly, from some initial value $\sigma_{1}$ to a $\sigma_{d-1}=c_{\sigma}\sigma_{1}$, $c_{\sigma}\in[0,1]$. To control spectral alignment, we set $h_{0}=\sum_{i=1}^{d-1}\alpha_{i}v_{i}$ subject to $\|\alpha\|_{\ell_{2}}=1$, and allow $\{\alpha_{i}\}_{i=1}^{d-1}$ to decay linearly from $\alpha_{1}$ to $\alpha_{d-1}=c_{\alpha}\alpha_{1}$, $c_{\alpha}\in[0,1]$. Each value of $c_{\alpha}\in[0,1]$ controls the alignment of $h_{0}$ with the top singular functions of $\mathcal{T}$: larger values correspond to stronger alignment. Similarly, $c_{\sigma}\in[0,1]$ controls the ill-posedness of the problem: smaller values indicate faster decay of singular values and hence more severe ill-posedness.
@@
In Figure˜1, we report the mean squared error (MSE) of two 2SLS estimators across a range of spectral decay rates ($c_{\sigma}$) and structural function decay rates ($c_{\alpha}$). Left: Oracle spectral features: The left panel uses 2SLS with fixed features $f$ and $g$ that form spectral features. This represents an oracle setting where feature learning is not needed. As expected, the MSE decreases from top to bottom as the problem becomes better conditioned (i.e., as $c_{\sigma}$ increases). Additionally, the MSE decreases from left to right, as smaller values of $c_{\alpha}$ correspond to stronger spectral alignment: more of $h_{0}$ is concentrated on the top singular functions of $\mathcal{T}$. These trends are fully consistent with the bound established in Corollary˜1. Right: Learned features. In the right panel, we learn 50 features from data by minimizing the empirical contrastive loss (8) using a neural network. We then build a 2SLS estimator on top of the learned features. The same qualitative trends are observed: MSE decreases with slower singular value decay and better alignment. Moreover, when $c_{\sigma}$ is small (top rows), learning the entire eigenspaces becomes difficult. In this regime, performance is most sensitive to $c_{\alpha}$. In particular, if $h_{0}$ is concentrated on the top singular functions (small $c_{\alpha}$), good recovery is still possible. In contrast, when $c_{\sigma}$ is large (bottom rows), the singular functions are equally important and more easily learned, rendering the alignment of $h_{0}$ less critical. The mismatch between the oracle and learned settings is because we do not exactly recover the singular feature spaces; as predicted by Theorem˜3. It is also worth noting that when $h_{0}$ is concentrated on high singular values, it may sometimes be beneficial for the singular values to decay quickly, by making it easier for the feature-learning model to learn the top singular functions well. This may explain the slight top-bottom increase in losses on the right end of the right panel.
The previous experiment assumes $h_{0}$ lies in the span of the singular functions of $\mathcal{T}$. Figure˜2 illustrates three regimes. The left panel shows a well-aligned case (“good/bad”), where $h_{0}$ is entirely supported on the singular functions of $\mathcal{T}$, allowing for exact recovery as long as enough data is available. The middle panel represents a “partial alignment” regime, where $h_{0}$ has some overlap with the singular functions; in this case, recovery is imperfect, but the estimator can still capture some signal. Finally, the right panel illustrates the “misaligned” or “very ugly” regime, where $h_{0}$ is orthogonal to the singular functions. The instrument provides no information about the signal, and all spectral methods fail. We refer to Appendix C for further discussion of these scenarios and the experimental setup.
@@ A.1 Proofs of Section˜2
Under Assumption˜2, the joint distribution $\pi_{X,Z}$ is absolutely continuous with respect to the product measure $\pi_{X}\otimes\pi_{Z}$, and thus admits a density
@@
In order to prove ˜1, we use the following lemma.
@@
Under Assumption˜2, for $\pi_{X}\otimes\pi_{Z}$-almost every $(x,z)$,
@@
By Lemma˜1, under Assumption˜2, for all $h\in L_{2}(X)$ and $\pi_{Z}\text{-a.e.}$,
$\mathcal{T}h(z)=\int_{\mathcal{X}}h(x)\,p(z,dx)=\int_{\mathcal{X}}h(x)p(x,z)d\pi_{X}(x)=\langle h,p(\cdot,z)\rangle_{L_{2}(X)}.$ (13)
We recall and prove ˜1. This is a classical result, see, e.g., [25].
@@
$=\sum_{i\geq 1}\int_{\mathcal{Z}}\left|\mathcal{T}e_{i}(z)\right|^{2}\,d\pi_{Z}(z)$ (expand $L_{2}(Z)$ norm)
$=\sum_{i\geq 1}\int_{\mathcal{Z}}\langle e_{i},p(\cdot,z)\rangle_{L_{2}(X)}^{2}\,d\pi_{Z}(z)$ (by Eq.˜13)
@@
$=\int_{\mathcal{Z}}\|p(\cdot,z)\|_{L_{2}(X)}^{2}\,d\pi_{Z}(z)$ (Parseval’s identity)
$=\int_{\mathcal{X}\times\mathcal{Z}}p(x,z)^{2}\,d\pi_{X}(x)\,d\pi_{Z}(z)$ (expand $L_{2}$ norm)
$<+\infty$ (by Assumption˜2, since $p\in L_{2}(\pi_{X}\otimes\pi_{Z})$).
@@ A.2 Proofs of Section˜3
˜4 is due to Lemma 1 of [3] and ˜2 corresponds to part (iii). We provide a full proof for completeness.
@@
We recall and prove Corollary˜1:
@@
Let us verify that Assumption˜3 holds under Definition˜2, i.e., when $\mathcal{H}_{\varphi,d}=\mathcal{V}_{d}$ and $\mathcal{H}_{\psi,d}=\mathcal{U}_{d}$.
@@
(ii) Let $h_{0}\in L_{2}(X)$, and consider its projection onto the orthogonal complement of $\mathcal{H}_{\varphi,d}$. Since $\mathcal{T}$ is compact with singular value decomposition $\mathcal{T}=\sum_{i=1}^{\infty}\sigma_{i}u_{i}\otimes v_{i}$ under Assumption˜2, we have
@@
This implies, using that $\tau_{\varphi,d}=\sigma_{d}^{-1}$ by ˜2,
@@
Since both parts of Assumption˜3 hold, and $\tau_{\varphi,d}=\sigma_{d}^{-1}$ by ˜2, we may apply Theorem˜1, which yields the desired result. ∎
A.3 Proofs of Section˜4
The following result shows that Eq.˜9 holds if and only if $\operatorname{rank}(\mathcal{T})\leq d$.
@@ A.3 Proofs of Section˜4
$\Rightarrow$ Assume $p(x,z)=\varphi(x)^{\mathsf{\scriptscriptstyle T}}\psi(z)$. Then, by Eq.˜10, for all $h\in L_{2}(X)$,
$\mathcal{T}h(z)$ $=\int h(x)p(z,dx)=\int h(x)p(x,z)\,d\pi_{X}(x)$ (by Lemma˜1)
$=\int h(x)\varphi(x)^{\mathsf{\scriptscriptstyle T}}\psi(z)\,d\pi_{X}(x)=\psi(z)^{\mathsf{\scriptscriptstyle T}}\int h(x)\varphi(x)\,d\pi_{X}(x)$
@@
$\int f(x,z)\,$ $\varphi(x)^{\mathsf{\scriptscriptstyle T}}\psi(z)\,d\pi_{X}(x)d\pi_{Z}(z)=\sum_{i=1}^{d}\int f(x,z)\varphi_{i}(x)\psi_{i}(z)\,d\pi_{X}(x)d\pi_{Z}(z)$
$=\sum_{i=1}^{d}\int_{\mathcal{Z}}\psi_{i}(z)\left(\int_{\mathcal{X}}f(x,z)\varphi_{i}(x)\,d\pi_{X}(x)\right)d\pi_{Z}(z)$
$=\int_{\mathcal{Z}}\left(\sum_{i=1}^{d}\psi_{i}(z)\langle f(\cdot,z),\varphi_{i}\rangle\right)d\pi_{Z}(z)$
$=\int_{\mathcal{Z}}\left[\mathcal{T}f(\cdot,z)\right](z)\,d\pi_{Z}(z)\quad\text{(since $\mathcal{T}=\sum\psi_{i}\otimes\varphi_{i}$)}$
@@
In order to prove Theorem˜3 we prove intermediate results.
@@
Assumption˜5 holds if and only if $\operatorname{rank}(\hat{\mathcal{T}}_{d})=d$.
@@
Assume that Assumption˜5 holds, i.e., $\{\hat{\varphi}_{1},\ldots,\hat{\varphi}_{d}\}$ and $\{\hat{\psi}_{1},\ldots,\hat{\psi}_{d}\}$ are linearly independent. We show that $\operatorname{rank}(\hat{\mathcal{T}}_{d})=d$.
@@
Under Assumption˜5, the family $\{\hat{\varphi}_{i}\}$ is linearly independent, so $\Phi$ is surjective. Similarly, the family $\{\hat{\psi}_{i}\}$ is linearly independent, so $A$ is injective. Therefore, the composition $A\circ\Phi$ has image of dimension
@@
Under Assumption˜5, we can therefore write $\hat{\mathcal{T}}_{d}$ in the following SVD form:
@@
Under Assumption˜5, $\mathcal{H}_{\hat{\varphi},d}=\operatorname{span}\left\{\hat{v}_{1},\ldots,\hat{v}_{d}\right\}$ and $\mathcal{H}_{\hat{\psi},d}=\operatorname{span}\left\{\hat{u}_{1},\ldots,\hat{u}_{d}\right\}$.
@@
We now recall and prove Theorem˜3: See 3
@@
We first show that $\sigma_{d}^{-1}\leq\tau_{\hat{\varphi},d}\leq(\sigma_{d}-2\hat{\varepsilon}_{d})^{-1}$. First note that if $h\in\mathcal{H}_{\hat{\varphi},d}$, then by Proposition 7, $h\in\mathcal{V}_{d}$. Using the SVD of $\hat{\mathcal{T}}_{d}$ from Eq.˜14, and writing $h=\sum_{i=1}^{d}\langle h,\hat{v}_{i}\rangle\hat{v}_{i}$, we obtain:
@@
[…]
@@
To conclude we now check Assumption˜3. Let us introduce
@@
The unique use of Assumption˜3-i) in the proof of Theorem B.1 [6] is to prove that $s_{\hat{\psi},\hat{\varphi},d}^{-1}=O(\tau_{\hat{\varphi},d})$ as $d\to+\infty$, which we now directly prove. First, note that for any $h\in L_{2}(X),$
@@
[…]
@@
which is satisfied if $\hat{\varepsilon}_{d}=o(\sigma_{d}^{2})$. We therefore conclude that if $\hat{\varepsilon}_{d}=o(\sigma_{d}^{2})$ then $s_{\hat{\psi},\hat{\varphi},d}^{-1}=O(\tau_{\hat{\varphi},d})$. Finally, we check Assumption˜3-ii). First note that by definition of $\Pi_{\hat{\varphi},d}$ and $\hat{\mathcal{T}}_{d}$, we have $\hat{\mathcal{T}}_{d}(\Pi_{\hat{\varphi},d})_{\perp}=0$, therefore,
@@
We now prove the equivalence between Eq.˜6 and Eq.˜7. While the proof strategy follows that of Kostic et al. [24, Theorem 1], we include our own version for completeness, as the parametrization of the learned operator differs. Specifically, [24] directly approximate the truncated operator $\mathcal{T}_{d}$ using a singular value decomposition of the form $\mathbbm{1}_{\mathcal{Z}}\otimes\mathbbm{1}_{\mathcal{X}}+\sum_{i=2}^{d}\hat{\sigma}_{i}\hat{\psi}_{i}\otimes\hat{\varphi}_{i}$, where the singular values $\hat{\sigma}_{i}$ are learned explicitly via a separate network. In contrast, our approach learns only the feature maps $(\varphi_{i},\psi_{i})$ and defines the approximation $\mathcal{T}_{d}(\hat{\varphi},\hat{\psi})=\sum_{i=1}^{d}\hat{\psi}_{i}\otimes\hat{\varphi}_{i}$. As such, we provide a short self-contained derivation adapted to our setting for clarity. Recall that the Eckart–Young–Mirsky formulation of the objective is defined as